OpenStudy (chazz16):

Need Help Will Multiplying Monomials and Binomials. MEDAL WILL BE AWARDED. First Question Below ↓↓↓↓↓↓↓↓↓↓↓↓↓↓↓↓ -20x(x-5)

OpenStudy (bookworm14):

First you use the distributive property and distibute the -20x --> -20x * x -20x * -5 you you multiply the coefficients (numbers) and add variables (letters) -20x * x= -20x^2 -20x * -5 = 100x FINAL ANSWER: -20x^2 + 100 Did this help? :)

OpenStudy (chazz16):

-20x^2+100x Its Wrong Its Said This: NO.Multiply each term in the parentheses by the monomial. If the variables are the same,Multiply them by adding the exponents.

OpenStudy (bookworm14):

That is weird, that is EXACTLY what I just did, hmm, gimme a sec. What is this on?

OpenStudy (bookworm14):

I am 100% sure that is the correct answer. I just asked my teacher. (I go to school online )

OpenStudy (chazz16):

is there an x after 100

OpenStudy (chazz16):

@Bookworm14 did you multiply the negatives????

OpenStudy (chazz16):

The Answer Is -20x^2+300x

OpenStudy (bookworm14):

Yes there is an X after the 100 x_x I forgot to type it in omg im so sorry. and there is NO WAY that the answer is -20x^2+300x ?! how did they get 300????

OpenStudy (chazz16):

I Didn't I'm Doing Plato So When I Summited My Answer It Said It Was Wrong And Than Said The Answer Is -20x^2+300x

OpenStudy (bookworm14):

That is so strange! Maybe it was a technical error? I know I was on Khan Academy (its like Plato and Study Island) and there was an error I had to report. Do you have a teacher y ou can ask?

OpenStudy (chazz16):

This is How Plato Got The Answer:Multiply each term in parentheses by the Monomial: -20x(x-15)= -20x•X -20x • (-15) = 20x^2 + 300x And I Don't At The Moment.

OpenStudy (bookworm14):

so was -20x(x-5) the problem they gave to work out?

OpenStudy (chazz16):

This Is The Question

OpenStudy (bookworm14):

I don't know darlin, that just has to be an error lol. Im sorry you got it wrong! :(

OpenStudy (chazz16):

Your Correct Its A Computer Error.

OpenStudy (bookworm14):

yayy!! lol I knew I didn't do it wrong! did my explantion help you out at all?